ganeshie8 (ganeshie8):

show that 12345678987654321 ≡ 0 (mod 12345679)

OpenStudy (usukidoll):

so we're in modulo 12345679 ... and we need a zero remainder.. oh wow. x.x

OpenStudy (usukidoll):

like how many cycles of 12345679 do we have to go through to reach 12345678987654321

ganeshie8 (ganeshie8):

Yes im sure it has a pretty neat solution :)

imqwerty (imqwerty):

i've done this problem before :D

Parth (parthkohli):

Facebook pages have taught me that \(111111111^2 = 12345678987654321\)

OpenStudy (usukidoll):

-_-! This is from the same book that I've used last year. I'm not sure if I made my professor solve this one lol

ganeshie8 (ganeshie8):

Haha what has that anything to do with the present problem

Parth (parthkohli):

The missing 8 in 12345679 is mildly annoying.

OpenStudy (usukidoll):

that's number 7b. page 52 x)

OpenStudy (usukidoll):

personally , I like the previous problems on page 51 XD

Parth (parthkohli):

\[12345679 \times 10^9 = 12345679000000000\]subtract 12345679 from this number. Woohoo.

Parth (parthkohli):

I know that it's ugly, but what's a better way to show that something divides something other than actually finding their ratio? :P

OpenStudy (anonymous):

12345678987654321=12345679987654321-1000000000 =12345679000000000+987654321-1000000000 the first term divided by 12345679 has remainder =0 And -1000000000+987654321=-12345679 divided by 12345679 remain 0 hence 12345678987654321\(\equiv\) 0 (mod 12345679)

OpenStudy (zzr0ck3r):

12345678987654321=999999999*12345679

OpenStudy (zzr0ck3r):

qed

OpenStudy (ikram002p):

was thinking of that the moment i saw it @zzr0ck3r :P

ganeshie8 (ganeshie8):

that factorization is pretty @ParthKohli / @OOOPS method is really clever!